OpenStudy (anonymous):

it takes Ron 12 hours to paint a house and it takes Rita 6 hours. How long should it take if they worked together? A. None of these choices are correct B. 18 hours C. 3 hours D. 9 hours E. 4 hours

OpenStudy (faman39):

18

OpenStudy (anonymous):

nope 18 only makes it longer. it should be shorter since they work together lol

OpenStudy (anonymous):

9hours

OpenStudy (anonymous):

(12+6)/2

OpenStudy (anonymous):

but wait.....working together should take at most a shorter time than the fastest worker...

OpenStudy (anonymous):

So I'm gonna stick with 4hours

OpenStudy (anonymous):

@Dyiliq are you sure it's for a test and i don't want to fail

OpenStudy (anonymous):

I couldn't give you a formula, but that seems reasonable... honestly, if she can do it in half the time he can do it, then both of them should be able to cut her time in half yet again so maybe its 3 hours

OpenStudy (anonymous):

it's 4 hours!!!!!!!!!!!11

OpenStudy (anonymous):

yeah, ultimate answer: 3hours

OpenStudy (anonymous):

4 hours and here's why:

OpenStudy (anonymous):

i'm so confused now

OpenStudy (anonymous):

If Ron can paint a house in 12 hours, then he can paint 1/12 of a house in 1 hour. If Rita can paint a house in 6 hours, then he can paint 1/6 of a house in 1 hour. Let x be the number of hours it would take them to paint the house. Working together, they can paint 1/x of the house in one hour, so 1/12+ 1/6= 1/x 3/12= 1/x Solve fpr LCD 3x=12 x=4

OpenStudy (anonymous):

@herschelle you're so smart

OpenStudy (anonymous):

my bad i spelled that wrong

OpenStudy (anonymous):

With a few help, of course. I didn't do all of it myself especially the verbosity (I'm kind of lazy) rest assured the answer is 4 Welcome!